# MESTI-MEGANet: Micro-expression Spatio-Temporal Image and Micro-expression Gradient Attention Networks for Micro-expression Recognition

### v1

@@ Abstract
Micro-expression recognition (MER) is a challenging task due to the subtle and fleeting nature of micro-expressions. Traditional input modalities, such as Apex Frame, Optical Flow, and Dynamic Image, often fail to adequately capture these brief facial movements, resulting in suboptimal performance. In this study, we introduce the Micro-expression Spatio-Temporal Image (MESTI), a novel dynamic input modality that transforms a video sequence into a single image while preserving the essential characteristics of micro-movements. Additionally, we present the Micro-expression Gradient Attention Network (MEGANet), which incorporates a novel Gradient Attention block to enhance the extraction of fine-grained motion features from micro-expressions. By combining MESTI and MEGANet, we aim to establish a more effective approach to MER. Extensive experiments were conducted to evaluate the effectiveness of MESTI, comparing it with existing input modalities across three CNN architectures (VGG19, ResNet50, and EfficientNetB0). Moreover, we demonstrate that replacing the input of previously published MER networks with MESTI leads to consistent performance improvements. The performance of MEGANet, both with MESTI and Dynamic Image, is also evaluated, showing that our proposed network achieves state-of-the-art results on the CASMEII and SAMM datasets. The combination of MEGANet and MESTI achieves the highest accuracy reported to date, setting a new benchmark for micro-expression recognition. These findings underscore the potential of MESTI as a superior input modality and MEGANet as an advanced recognition network, paving the way for more effective MER systems in a variety of applications.
@@
Micro-expresison Spatio-Temporal Image, gradient attention, micro-expression, micro-expression recognition, micro-expression input modality, micro-expression recognition network.
@@ I INTRODUCTION
Facial expression, a vital channel of non-verbal communication, encompasses two primary types: macro expressions and micro expressions. Macro expressions are typically deliberate, easily observable, and last for an extended period, conveying a person’s emotions openly [1]. In contrast, micro expressions (ME) are brief, involuntary facial movements that last less than 0.5 seconds [3, 2], making them significantly challenging to control or fabricate. Unlike macro expressions, micro expressions reveal a person’s genuine emotions, often surfacing when one attempts to conceal their true feelings [4]. These fleeting expressions are especially revealing in high-risk situations [5, 6], where concealing emotions is common. Since these unique characteristics of micro expressions, they have garnered significant attention as a channel for uncovering individuals’ genuine thoughts and emotions. Their involuntary nature provides valuable insights, making them highly applicable in a range of critical fields. For instance, micro-expressions play a crucial role in enhancing the accuracy of deception detection systems, providing valuable insights that more prolonged expressions may not capture [7]. In criminal investigations, law enforcement officers can assess a suspect’s truthfulness by analyzing micro-expressions that may contradict verbal statements [8]. Beyond security applications, micro-expressions are increasingly relevant in healthcare, particularly in clinical settings, where they can provide essential clues about a patient’s emotional state and aid medical professionals in assessing recovery progress [9].
Despite its potential, micro-expression recognition (MER) presents significant challenges due to the brevity and subtle intensity of micro-expressions. Studies have shown that even experts achieve only 47% accuracy in recognizing micro-expressions, highlighting the inherent complexity of this task [10]. However, leveraging advancements in computational capabilities, as well as modern machine learning and deep learning algorithms, computer-based systems for micro-expression analysis have demonstrated significant superiority over human performance, with accuracy rates often exceeding 50%. These advancements offer a promising pathway for achieving more accurate and reliable recognition of micro-expressions across a wide range of applications [11].
@@
Static Images: Static image-based approaches often utilize the apex frame, which represents the peak intensity of a micro-expression. This method reduces the complexity of video data by condensing an entire sequence into a single frame, thereby decreasing computational overhead. However, static images lack temporal information, which is crucial for identifying micro-expressions characterized by their rapid onset and offset. The absence of these temporal cues may lead to misclassification, as the system cannot discern the subtle temporal changes critical for distinguishing micro-expressions. Consequently, static image-based MER systems often exhibit suboptimal accuracy [13, 12].
@@
Optical Flow: Optical flow, a widely used technique for motion representation, provides both the magnitude and direction of pixel movement through a two-dimensional vector field comprising horizontal and vertical flows. This method introduces temporal information into the analysis, enhancing the system’s ability to capture micro-expression dynamics. However, the effective integration of optical flow in MER systems often requires the design of multi-stream network architectures to process both horizontal and vertical flow components. Additionally, optical flow techniques are susceptible to noise caused by lighting variations, head movements, and other extraneous factors, potentially impairing recognition accuracy. [17, 27]
@@
Dynamic Imaging Methods: Dynamic imaging methods summarize a video sequence into a single image that encodes temporal information. A notable example of this technique is the Dynamic Image [28], which represents a video as a single image capturing the overall temporal essence of the sequence. This method has been shown to be effective in action recognition tasks, as demonstrated by the original authors. Subsequently, it has been adapted for micro-expression recognition (MER) in several studies [15, 56]. However, the direct inheritance of this technique from action recognition tasks has limited its effectiveness in MER, as it struggles to accurately capture the subtle and minute facial movements characteristic of micro-expressions. Recognizing the inadequacy of existing input representations for MER, some researchers have attempted to refine dynamic imaging techniques specifically for this domain. For instance, Affective Image [32] and Active Image [16] are examples of efforts to tailor dynamic imaging approaches to better suit MER tasks. While these studies made progress in adapting the dynamic image concept, their input representations remain limited in comprehensiveness. Both approaches designed specialized networks for their respective inputs, yet the recognition accuracy of these networks has remained constrained, achieving only 50–60% on four-class classification tasks, rather than the five-class standard used in earlier research.
@@
In summary, existing input modalities fail to comprehensively address the challenges of micro-expression recognition. Static images, while computationally efficient, lack temporal information critical for capturing transient micro-movements. Optical flow, though effective in encoding motion, struggles to represent the subtle intensity and brevity of micro-expressions, often resulting in noisy estimations due to their low-amplitude characteristics. These limitations underscore the pressing need for a modality that seamlessly integrates nuanced spatio-temporal motion cues into a compact and discriminative representation, tailored specifically for micro-expression dynamics.
@@
On the other hand, although Convolutional Neural Networks (CNN) have demonstrated effectiveness in extracting spatial features from facial expressions, they face notable limitations when applied to MER. Traditional CNN architectures tend to focus on features with prominent magnitudes, often overlooking the extremely subtle and transient motion signals inherent in micro-expressions. This underrepresentation of fine-grained motion details reduces the model’s ability to distinguish between different micro-expressions, particularly when the movements are brief and of low intensity. These challenges highlight the need for more specialized architectures and input representations that can effectively capture and amplify subtle motion cues.
@@
Previous studies have proposed various network architectures to address these limitations, yet significant gaps remain. For instance, MER-GCN [53] employs Graph Convolutional Networks (GCNs) to model the spatio-temporal relationships between Action Units (AUs). While this approach shows promise, it suffers from high computational complexity and limited generalization across diverse datasets such as SAMM. Additionally, the use of AUs as input modality is suboptimal, as accurately identifying AUs in real-world conditions is still highly challenging. LEARNet [51] utilizes Dynamic Image as input with a standard CNN architecture but lacks a dedicated attention mechanism tailored for micro-movements, leading to the omission of critical features. GEME [56] introduces a multi-task framework that incorporates gender information, but its feature extraction process still fails to fully leverage subtle gradient variations. AMAN [52] integrates an attention mechanism into CNNs to focus on facial regions; however, this mechanism primarily relies on raw pixel intensity rather than gradient-based motion, making it less effective in detecting low-intensity changes. Similarly, CapsNet [50] and optical flow-based methods using OFF-ApexNet [46] have shown improved accuracy but remain susceptible to noise caused by lighting variations and head movements, particularly for micro-expressions with low amplitude.
@@
These limitations underscore the need for a more robust and efficient approach that can amplify subtle motion informations and direct attention to regions exhibiting significant gradient changes. To address these challenges, in this paper, a novel MER method is proposed that introduces a dynamic input modality called the Micro-expression Spatio-Temporal Image (MESTI). This modality represents a video sequence as a single image by capturing the motion features of micro-expressions. Furthermore, this paper introduces introduce Micro-expression Gradient Attention Networks (MEGANet), a novel network architecture that incorporates a Gradient Attention mechanism. This mechanism leverages gradient information from intermediate layers to highlight regions with subtle motion, thereby enhancing the extraction of fine-grained features that are critical for accurate micro-expression recognition. The contributions of the paper are as follows:
@@
  • A novel input modality, MESTI, which represents video sequences as single images that specifically synthesize and highlight micro-movements within micro-expression videos.
@@
  • MEGANet: By integrating a novel Gradient Attention Block, we develop a micro-expression network capable of focusing on motion regions, thereby improving the performance of micro-expression recognition.
@@
Through extensive experiments, the effectiveness of each proposed component (MESTI, MEGANet) is demonstrated by evaluating their individual contributions and their combination with previously published methods. The results show that each component of our proposed method enhances the performance of the micro-expression recognition (MER) process, and when combined, MESTI and MEGANet yield a effective overall MER approach.
@@ II-A Micro-expression Spatio-Temporal Image
The initial idea for creating an effective input representation for ME stemmed from observing and studying the motion characteristics of micro-expressions. The intensity of motion gradually increases from the onset frame (the starting frame) to the apex (the frame with the highest ME intensity), then decreases towards the offset frame (the final frame representing the ME). Based on this characteristic, the proposed method simulates this motion in the process of constructing a distinctive representation for micro-expressions, namely MESTI. Our objective is to create a spatio-temporal image that effectively represents a micro-expression video. To achieve this, a temporal encoding approach introduced that transforms the entire video sequence into a single representative image. Additionally, our method incorporates the process of aggregating spatial information from the video into a compact static representation.
Inspired by the approximate rank pooling method, which has been used in modeling video evolution [29], a similar strategy is proposed to encode the temporal evolution of micro-expressions into a single image. This approach captures the dynamic variations in facial expressions over time while preserving the spatial structure necessary for effective micro-expression recognition.
@@ II-A1 Spatial Encoding
This process integrates spatial information from individual frames into a micro-expression image that preserves structural and appearance details. By leveraging the extracted RGB feature vectors, the method ensures that spatial characteristics of each frame are considered in the ranking process, allowing the network to learn an optimal frame-ordering that reflects their relative importance in the sequence.
@@ II-A2 Temporal Encoding
Temporal encoding is performed based on the characteristic motion patterns of micro-expressions, which serve as a basis for assigning scores to each frame during the rank pooling process of spatial encoding. Figure 1 illustrates the intensity of motion in ME. The motion characteristics of micro-expressions can be easily observed: the intensity gradually increases from the first frame (onset), peaks at the apex frame, and then gradually decreases toward the final frame (offset) of the micro-expression. Therefore, in this study, we aim to model the motion characteristics of micro-expressions within the temporal encoding process to construct a micro-expression image from the video.
Temporal encoding is implemented by generating a ranking score that simulates the motion intensity of the micro-expression in a straightforward manner during the rank pooling process. Let $I_{a}$ defined as the apex frame, where the motion intensity of the micro-expression reaches its maximum. Given any two frames $I_{q},I_{t}$, the frame closer to the apex frame is assigned a higher ranking score in our ranking function.
@@
The MESTI construction result is shown in Figure 2b using frame sequence (Figure 2a) as input. From the input frame sequence, we represent and observe the motion intensity of the ME representation and have the graph below. The MESTI construction results show that, firstly, our method generates a ranking function that better simulates the nature of the micro-expression motion. Second, through the visual representation results, MESTI has shown more clearly the action units in the micro-expression on the final image constructed compared to the traditional dynamic image method as showed in Figure 2c.
Figure 3: The architecture of the proposed Micro-Expression Gradient Attention Network (MEGANet)
@@ II-B Micro-expression Gradient Attention Networks
The challenge in micro-expression recognition lies in capturing the subtle, transient spatiotemporal patterns that characterize micro-expressions, which often involve subtle intensity changes that conventional CNNs struggle to detect. These expressions are fleeting, making it difficult for traditional methods to effectively focus on the most critical regions of motion. To address this, MEGANet is proposed, a MER network that aims to enhance the detection of micro-expressions by directing attention to areas with significant gradient changes. The core idea behind MEGANet is to combine gradient-guided attention with spatial self-attention, enabling the network to focus on both fine-grained motion transitions and the broader spatial context.
@@
  • Gradient Attention Block: Computes spatial gradients to enhance subtle micro-expression features. A convolutional layer followed by a sigmoid activation generates an attention map, which is multiplied with the original input to highlight key regions.
@@
This architecture effectively captures micro-expression dynamics by leveraging gradient-based attention and residual learning, improving the network’s ability to recognize subtle facial movements.

### v2

@@ Abstract
Micro-expression recognition (MER) is a challenging task due to the subtle and fleeting nature of micro-expressions. Traditional input modalities, such as Apex Frame, Optical Flow, and Dynamic Image, often fail to adequately capture these brief facial movements, resulting in suboptimal performance. In this study, we introduce the Micro-expression Spatio-Temporal Image (MESTI), a novel dynamic input modality that transforms a video sequence into a single image while preserving the essential characteristics of micro-movements. Additionally, we present the Micro-expression Gradient Attention Network (MEGANet), which incorporates a novel Gradient Attention block to enhance the extraction of fine-grained motion features from micro-expressions. By combining MESTI and MEGANet, we aim to establish a more effective approach to MER. Extensive experiments were conducted to evaluate the effectiveness of MESTI, comparing it with existing input modalities across regular architectures. Moreover, we demonstrate that replacing the input of previously published MER networks with MESTI leads to consistent performance improvements. The performance of MEGANet is also evaluated, showing that our proposed network achieves state-of-the-art results on the SMIC-HS, SAMM and competitive performance on CASMEII datasets. The combination of MEGANet and MESTI achieves the highest accuracy reported to date, setting a new benchmark for micro-expression recognition. These findings underscore the potential of MESTI as a superior input modality and MEGANet as an advanced recognition network, aiming to more effective MER systems in a variety of applications.
@@
Micro-expression, micro-expression recognition, gradient attention, micro-expression input modality, micro-expression recognition network.
@@ I INTRODUCTION
Facial expression, a vital channel of non-verbal communication, encompasses two primary types: macro expressions and micro expressions. Macro expressions are typically deliberate, easily observable, and last for an extended period, conveying a person’s emotions openly [1]. In contrast, micro-expressions (MEs) are brief, involuntary facial movements that last less than 0.5 seconds [3, 2], making them significantly challenging to control or fabricate. Unlike macro expressions, MEs reveal a person’s genuine emotions, often surfacing when one attempts to conceal their true feelings [4]. These fleeting expressions are especially revealing in high-risk situations [5, 6], where concealing emotions is common. Since these unique characteristics of MEs, they have garnered significant attention as a channel for uncovering individuals’ genuine thoughts and emotions. Their involuntary nature provides valuable insights, making them highly applicable in a range of critical fields. For instance, MEs play a crucial role in enhancing the accuracy of deception detection systems, providing valuable insights that more prolonged expressions may not capture [7]. In criminal investigations, law enforcement officers can assess a suspect’s truthfulness by analyzing MEs that may contradict verbal statements [8]. Beyond security applications, MEs are increasingly relevant in healthcare, particularly in clinical settings, where they can provide essential clues about a patient’s emotional state and aid medical professionals in assessing recovery progress [9].
Despite its potential, ME recognition (MER) presents significant challenges due to the brevity and subtle intensity of MEs. Studies have shown that even experts achieve only 47% accuracy in recognizing MEs, highlighting the inherent complexity of this task [12]. However, leveraging advancements in computational capabilities, as well as modern machine learning and deep learning algorithms, computer-based systems for ME analysis have demonstrated significant superiority over human performance, with accuracy rates often exceeding 50%. These advancements offer a promising pathway for achieving more accurate and reliable recognition of MEs across a wide range of applications [13].
@@
Static Images: Static image-based approaches often utilize the apex frame, which represents the peak intensity of a ME. This method reduces the complexity of video data by condensing an entire sequence into a single frame, thereby decreasing computational overhead. However, static images lack temporal information, which is crucial for identifying MEs characterized by their rapid onset and offset. The absence of these temporal cues may lead to misclassification, as the system cannot discern the subtle temporal changes critical for distinguishing MEs. Consequently, static image-based MER systems often exhibit suboptimal accuracy [15, 14].
@@
Optical Flow: Optical flow, a widely used technique for motion representation, provides both the magnitude and direction of pixel movement through a two-dimensional vector field comprising horizontal and vertical flows. This method introduces temporal information into the analysis, enhancing the system’s ability to capture ME dynamics. However, the effective integration of optical flow in MER systems often requires the design of multi-stream network architectures to process both horizontal and vertical flow components. Additionally, optical flow techniques are susceptible to noise caused by lighting variations, head movements, and other extraneous factors, potentially impairing recognition accuracy. [19, 34]
@@
Dynamic Imaging Methods: Dynamic imaging methods summarize a video sequence into a single image that encodes temporal information. A notable example of this technique is the Dynamic Image [36], which represents a video as a single image capturing the overall temporal essence of the sequence. This method has been shown to be effective in action recognition tasks, as demonstrated by the original authors. Subsequently, it has been adapted for ME recognition in several studies [17, 68]. However, the direct inheritance of this technique from action recognition tasks has limited its effectiveness in MER, as it struggles to accurately capture the subtle and minute facial movements characteristic of MEs. Recognizing the inadequacy of existing input representations for MER, some researchers have attempted to refine dynamic imaging techniques specifically for this domain. For instance, Affective Image [41] and Active Image [18] are examples of efforts to tailor dynamic imaging approaches to better suit MER tasks. While these studies made progress in adapting the dynamic image concept, their input representations remain limited in comprehensiveness. Both approaches designed specialized networks for their respective inputs, yet the recognition accuracy of these networks has remained constrained, achieving only 50–60% on four-class classification tasks, rather than the five-class standard used in earlier research.
@@
In summary, existing input modalities fail to comprehensively address the challenges of ME recognition. Static images, while computationally efficient, lack temporal information critical for capturing transient micro-movements. Optical flow, though effective in encoding motion, struggles to represent the subtle intensity and brevity of MEs, often resulting in noisy estimations due to their low-amplitude characteristics. These limitations underscore the pressing need for a modality that seamlessly integrates nuanced spatio-temporal motion cues into a compact and discriminative representation, tailored specifically for ME dynamics.
@@
On the other hand, although Convolutional Neural Networks (CNN) have demonstrated effectiveness in extracting spatial features from facial expressions, they face notable limitations when applied to MER. Traditional CNN architectures tend to focus on features with prominent magnitudes, often overlooking the extremely subtle and transient motion signals inherent in MEs. This underrepresentation of fine-grained motion details reduces the model’s ability to distinguish between different MEs, particularly when the movements are brief and of low intensity. These challenges highlight the need for more specialized architectures and input representations that can effectively capture and amplify subtle motion cues.
@@
Previous studies have proposed various network architectures to address these limitations, yet significant gaps remain. For instance, MER-GCN [66] employs Graph Convolutional Networks (GCNs) to model the spatio-temporal relationships between Action Units (AUs). While this approach shows promise, it suffers from high computational complexity and limited generalization across diverse datasets such as SAMM. Additionally, the use of AUs as input modality is suboptimal, as accurately identifying AUs in real-world conditions is still highly challenging. LEARNet [63] utilizes Dynamic Image as input with a standard CNN architecture but lacks a dedicated attention mechanism tailored for micro-movements, leading to the omission of critical features. GEME [68] introduces a multi-task framework that incorporates gender information, but its feature extraction process still fails to fully leverage subtle gradient variations. AMAN [64] integrates an attention mechanism into CNNs to focus on facial regions; however, this mechanism primarily relies on raw pixel intensity rather than gradient-based motion, making it less effective in detecting low-intensity changes. Similarly, CapsNet [62] and optical flow-based methods using OFF-ApexNet [54] have shown improved accuracy but remain susceptible to noise caused by lighting variations and head movements, particularly for MEs with low amplitude.
@@
These limitations underscore the need for a more robust and efficient approach that can amplify subtle motion informations and direct attention to regions exhibiting significant gradient changes. To address these challenges, in this paper, a novel MER method is proposed that introduces a dynamic input modality called the ME Spatio-Temporal Image (MESTI). This modality represents a video sequence as a single image by capturing the motion features of MEs. Furthermore, this paper introduces introduce ME Gradient Attention Networks (MEGANet), a novel network architecture that incorporates a Gradient Attention mechanism. This mechanism leverages gradient information from intermediate layers to highlight regions with subtle motion, thereby enhancing the extraction of fine-grained features that are critical for accurate ME recognition. The contributions of the paper are as follows:
@@
  • A novel input modality, MESTI, which represents video sequences as single images that specifically synthesize and highlight micro-movements within ME videos.
@@
  • MEGANet: By integrating a novel Gradient Attention Block and Residual Attention Block, we develop a ME network capable of focusing on motion regions, thereby improving the performance of ME recognition.
  • A comprehensive set of experimental scenarios is designed to validate the effectiveness of the proposed components, achieving performance that outperforms previous state-of-the-art studies.
@@
Through extensive experiments, the effectiveness of each proposed component (MESTI, MEGANet) is demonstrated by evaluating their individual contributions and their combination with previously published methods. The results show that each component of our proposed method enhances the performance of the ME recognition process, and when combined, MESTI and MEGANet yield a effective overall MER approach.
@@ II-A Micro-expression Spatio-Temporal Image
The initial idea for creating an effective input representation for ME stemmed from observing and studying the motion characteristics of MEs. The intensity of motion gradually increases from the onset frame (the starting frame) to the apex (the frame with the highest ME intensity), then decreases towards the offset frame (the final frame representing the ME). Based on this characteristic, the proposed method simulates this motion in the process of constructing a distinctive representation for MEs, namely MESTI. Our objective is to create a spatio-temporal image that effectively represents a ME video. To achieve this, a temporal encoding approach introduced that transforms the entire video sequence into a single representative image. Additionally, our method incorporates the process of aggregating spatial information from the video into a compact static representation.
Inspired by the approximate rank pooling method, which has been used in modeling video evolution [37], a similar strategy is proposed to encode the temporal evolution of MEs into a single image. This approach captures the dynamic variations in facial expressions over time while preserving the spatial structure necessary for effective ME recognition.
@@ II-A1 Spatial Encoding
This process integrates spatial information from individual frames into a ME image that preserves structural and appearance details. By leveraging the extracted RGB feature vectors, the method ensures that spatial characteristics of each frame are considered in the ranking process, allowing the network to learn an optimal frame-ordering that reflects their relative importance in the sequence.
@@ II-A2 Temporal Encoding
Temporal encoding is performed based on the characteristic motion patterns of MEs, which serve as a basis for assigning scores to each frame during the rank pooling process of spatial encoding. Figure 1 illustrates the intensity of motion in ME. The motion characteristics of MEs can be easily observed: the intensity gradually increases from the first frame (onset), peaks at the apex frame, and then gradually decreases toward the final frame (offset) of the ME. Therefore, in this study, we aim to model the motion characteristics of MEs within the temporal encoding process to construct a ME image from the video.
Temporal encoding is implemented by generating a ranking score that simulates the motion intensity of the ME in a straightforward manner during the rank pooling process. Let $I_{a}$ defined as the apex frame, where the motion intensity of the ME reaches its maximum. Given any two frames $I_{q},I_{t}$, the frame closer to the apex frame is assigned a higher ranking score in our ranking function.
@@
The MESTI construction result is shown in Figure 2b using frame sequence (Figure 2a) as input. From the input frame sequence, we represent and observe the motion intensity of the ME representation and have the graph below. The MESTI construction results show that, firstly, our method generates a ranking function that better simulates the nature of the ME motion. Second, through the visual representation results, MESTI has shown more clearly the action units in the ME on the final image constructed compared to the traditional dynamic image method as showed in Figure 2c.
Figure 3: The architecture of the proposed ME Gradient Attention Network (MEGANet)
@@ II-B Micro-expression Gradient Attention Networks
The challenge in ME recognition lies in capturing the subtle, transient spatiotemporal patterns that characterize MEs, which often involve subtle intensity changes that conventional CNNs struggle to detect. These expressions are fleeting, making it difficult for traditional methods to effectively focus on the most critical regions of motion. To address this, MEGANet is proposed, a MER network that aims to enhance the detection of MEs by directing attention to areas with significant gradient changes. The core idea behind MEGANet is to combine gradient-guided attention with spatial self-attention, enabling the network to focus on both fine-grained motion transitions and the broader spatial context.
@@
  • Gradient Attention Block: Computes spatial gradients to enhance subtle ME features. A convolutional layer followed by a sigmoid activation generates an attention map, which is multiplied with the original input to highlight key regions.
@@
This architecture effectively captures ME dynamics by leveraging gradient-based attention and residual learning, improving the network’s ability to recognize subtle facial movements.